# ThermoHands: A Benchmark for 3D Hand Pose Estimation from Egocentric Thermal Images

### v3

@@ 4.4.2 Quantitative Results under Normal Conditions
HTT (Sequence) TherFormer-V TherFormer-S Average
RGB 43.30 0.542 43.50 0.542 44.61 0.529 43.80 0.538
Depth 41.62 0.559 39.70 0.581 39.84 0.579 40.39 0.573
NIR 41.57 0.562 40.79 0.575 40.98 0.575 41.11 0.571
Thermal 47.07 0.512 43.36 0.549 44.64 0.539 45.06 0.533
Table 4: Comparison between different spectra under normal conditions. Models are trained with their corresponding spectrum images from the training set. We test them on the main testing set.
5 Limitation and Future Work
Despite the advancements, we acknowledge certain limitations here guiding our future work. First, the scale of our current dataset is relatively small when compared to some datasets [29, 40, 39, 38]. We plan to continuously enlarge this dataset by incorporating data from new scenarios and hand actions. Second, we only annotate 3D hand pose for our dataset, which limits its usage of evaluation to other tasks relevant to human hands. Further efforts will be annotating the fine-grained hand action splits and contact heatmap for hand grasp [42]. Lastly, we currently rely on qualitative analysis for challenging conditions when our automatic annotation method becomes unreliable. In the future, we will leverage unaffected spectral images in individual cases to manually annotate the 3D hand pose.
6 Conclusion
@@
This paper introduces ThermoHands, the first benchmark for egocentric 3D hand pose estimation using thermal images. It features a multi-spectral, multi-view dataset with automatically annotated 3D poses and a novel baseline method utilizing dual transformer modules for encoding spatio-temporal relationships. We demonstrate near 1cm annotation accuracy, show that TherFormer surpasses existing methods in thermal-based 3D hand pose estimation, and confirm thermal images’ effectiveness in challenging lighting and obstruction scenarios. We believe our endeavour could pave the way for further research in thermal-based 3D hand pose estimation and its wide application.
Acknowledgments and Disclosure of Funding
This research is partially supported by the Engineering and Physical Sciences Research Council (EPSRC) under the Centre for Doctoral Training in Robotics and Autonomous Systems at the Edinburgh Centre of Robotics (EP/S023208/1), and partially supported by Cisco Research.
Checklist
  1. 1.
    For all authors…
    1. (a)
      Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes] We elaborate our contributions as a list in the introduction section and briefly explain them in the abstract.
    2. (b)
      Did you describe the limitations of your work? [Yes] See Sec. 5
    3. (c)
      Did you discuss any potential negative societal impacts of your work? [N/A] There is no societal impact of this work.
    4. (d)
      Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes] We have reviewed the ethics review guidelines and make sure we obey all rules for this research.
  2. 2.
    If you are including theoretical results…
    1. (a)
      Did you state the full set of assumptions of all theoretical results? [N/A] This work does not contribute to the theoretical results.
    2. (b)
      Did you include complete proofs of all theoretical results? [N/A] This work does not contribute to the theoretical results.
  3. 3.
    If you ran experiments (e.g. for benchmarks)…
    1. (a)
      Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? [No] We will release our code upon the acceptance of this paper.
    2. (b)
      Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? [Yes] We describe them in our main paper and supplementary document.
    3. (c)
      Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)? [No] In line with the state-of-the-art 3D hand pose estimation works, we do not report error bars for our experiment results.
    4. (d)
      Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes] We include the type of GPUs in the main paper.
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
    1. (a)
      If your work uses existing assets, did you cite the creators? [Yes] We use the offical code of HTT [15] and A2J-transformer [46] to reproduce their results on our dataset, which have been cited.
    2. (b)
      Did you mention the license of the assets? [Yes] We mention the license of the offical codes of HTT and A2J-Transformer in the supplementary document.
    3. (c)
      Did you include any new assets either in the supplemental material or as a URL? [Yes] We include the datasheet and the download URL of our dataset in the supplemental document.
    4. (d)
      Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [Yes] We claim in the footnote that participant consent forms were signed before the collection.
    5. (e)
      Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [Yes] We discuss in the supplementary document.
  5. 5.
    If you used crowdsourcing or conducted research with human subjects…
    1. (a)
      Did you include the full text of instructions given to participants and screenshots, if applicable? [N/A] We do not give written instructions to our participants.
    2. (b)
      Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [N/A] There is no potential participant risks.
    3. (c)
      Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [N/A] Our participants are volunteer.

### v4

@@ 6.4.3. Quantitative Results under Normal Conditions
7. Conclusion
@@
This paper introduces ThermoHands, the first benchmark for egocentric 3D hand pose estimation using thermal images. ThermoHands features a multi-spectral, multi-view dataset with automatically annotated 3D hand poses and a novel baseline method, TherFormer, utilizing dual transformer modules for encoding spatio-temporal relationships. We demonstrate near 1cm annotation accuracy, show that TherFormer surpasses existing methods in thermal-based 3D hand pose estimation, and confirm thermal images’ effectiveness in challenging lighting and obstruction scenarios. We believe our foundational endeavour could set the stage for further research in thermal-based 3D hand pose estimation and its wide application.
Limitation. As the first exploration in this field, this work has certain limitations that can guide our future work. For example, we only annotate 3D hand pose for our dataset, which limits its usage of evaluation to other tasks relevant to human hands. Further efforts could be annotating the fine-grained hand action splits and hand-object contact (brahmbhatt2020contactpose, ).
Appendix
The appendix is organized as follows:
  • $\bullet$
    Appendix A illustrates more details about our hand pose dataset, in the aspects of spectrum coverage, interference avoidance, and dataset distribution.
  • $\bullet$
    Appendix B presents more implementation details of our automatic 3D hand pose annotation method.
  • $\bullet$
    Appendix C introduce more architecture design, parameter and loss details of TherFormer.
We also provide more visualization in the format of both images and videos for our data, annotation and results on our anonymous project page.
Appendix A Multi-spectral Hand Pose Dataset
Spectrum Coverage and Interference Avoidance. The specification of all sensor frames we collect are shown in Tab. 1. Note that we keep the faces of subjects outside of the FoV of the exocentric camera to avoid any personally identifiable information. The multi-spectral data encompasses imaging from the visible spectrum (400-700nm), near-infrared (NIR) spectrum (850 nm $\pm$ 10 nm), and the LWIR spectrum (8-14 $\mu$m). Images captured across different spectra contain unique information and serve varied purposes. For instance, RGB images offer semantic information, facilitating a deeper understanding of human-environment interaction. NIR lights can be actively emitted by our depth cameras to obtain the depth measurements via ToF or structured lighting. The LWIR frame, capturing temperature information, readily isolates uniform heat emitters like human hands. On the head-mounted sensor platform, the L515 LiDAR depth camera (IntelRSL515, ) emits NIR lasers at a wavelength of 860 nm, which falls outside the thermal camera’s receptive range (8-14 $\mu$m), thereby eliminating any potential interference between cameras on the HMSP. Conversely, the exocentric RGB-D camera (IntelRealSense2023D455, ) necessitates structured lighting employing NIR at a wavelength identical to the IR emitter on the L515 LiDAR camera (IntelRSL515, ). To prevent interference and image corruption, the exocentric RGB-D camera and the egocentric NIR LiDAR are strategically positioned outside each other’s receptive fields during data collection. In the office environment, random heat sources, e.g., servers and chargers, are strategically placed in the background to increase realism and introduce challenging factors into thermal images.
Resolution Fov FPS
Sensor frames Sensor Type Range Horizontal Vertical Range Horizontal Vertical
RGB (ego) Intel RS L515 (IntelRSL515, ) - 1280 720 - 69 42 30
NIR (ego) Intel RS L515 (IntelRSL515, ) - 640 480 - 70 55 30
Depth (ego) Intel RS L515 (IntelRSL515, ) ¡ 5mm @ 1m 640 480 0.25m to 9m 70 55 30
Thermal (ego) FLIR Boson 640 (FLIRBoson2024, ) $\leq$ 60 mK 640 512 - 95 - 8.5
RGB (exo) Intel RS D455 (IntelRealSense2023D455, ) - 1280 720 - 87 58 30
[…]
Table 6. The specification of sensor frames captured in data collection.
Figure 9. Distribution of data from the main part over hand actions.
Dataset Distribution. We show the data distribution of our main part over hand actions in Fig. 9. Among hand-object interaction actions, write with pen, write with pencil and read book have more frames than others due to the complexity of these actions, making them tend to last longer than others. As seven of our participants did not perform the hand-virtual actions, we have fewer frames from them than their hand-object counterparts. Please see our project page for more visualization of our collected data.
Appendix B Dataset Annotation
Acquisition of 2D Keypoint and Mask. We utilize the open-source MediaPipe Hands pipeline (mediapipe_hands_2020, ) to infer the 2D hand keypoints from both egocentric and exocentric RGB images. To improve the recall of hand detection, we modify the hyperparameters min_detection_confidence from the default 0.5 to 0.1. Sequential RGB images are fed into the MeidaPipe Hands pipeline to obtain the 2D keypoints for two hands. Particularly, we distinguish between two hands according to their relative locations on images. Given 2D keypoints estimated from the task-specific MeidiaPipe Hands model, we employ the versatile Segment-Anything Model (SAM) (li2020segmenting, ) to infer the 2D hand masks. To ensure the high quality of the generated 2D hand mask, we utilize the largest version of SAM, i.e., ViT-L SAM model and prompt it with both the 2D hand keypoints and the bounding boxes defined by them. As a result, we acquire 2D hand keypoints and masks of two hands for each frame.
3D Keypoint Triangulation. After inferring the 2D hand keypoints from two views, we can obtain the positions of 3D hand keypoints using triangulation. To implement this, we use the OpenCV function cv2.triangulatePoints (opencv_library, ).
3D Hand Point Cloud Generation. To generate 3D hand point cloud, we first index the hand pixels with the 2D hand mask on the depth image and then convert them into 3D points with the camera intrinsic parameters. Points generated from the exocentric view are transformed into the egocentric camera space. By merging points from two viewpoints, we can obtain a dense 3D hand point cloud.
Joint Angle Limitation. Following (hampali2020honnotate, ; kwon2021h2o, ; liu2022hoi4d, ), we optimize the MANO model in the joint angle space instead of the PCA space so that we can constrain the joint angles explicitly. As said in the main paper, the joint angles are limited by a set of empirical boundary values during optimization. Specifically, we use the same parameters of the upper and lower joint angle boundaries as (kwon2021h2o, ).
Shape Regularization. To avoid unrealistic hand shape, we also add regularization to the shape parameter $\beta$ when optimizing it for the first frame of each sequence, which can be written as:
(7) $\mathcal{L}_{shape}=\sum_{i=1}^{10}||\beta_{i}||$
MANO Fitting. We utilize the Pytorch version of the MANO layer proposed in (hasson19_obman, ). The Adam optimizer (kingma2014adam, ) is used to minimize the optimization objective mentioned in the main paper. To obtain accurate 3D hand pose annotation, we optimize the MANO parameters individually for each frame instead of considering batches of them. For the first frame of each sequence, we run the optimization for 500 iterations with an initial learning rate of as 0.1 decayed by 0.9 for each 50 frames. For the subsequent frames, as we initialize with the results from the last frame, we only optimize for 60 iterations and initialize the learning rate as 0.05 to speed up the convergence.
Appendix C Baseline Method
Backbone. The input to our backbone is consecutive thermal images $\mathcal{S}=\{\mathcal{I}_{i}\}_{i=1}^{T}$ resized to 320$\times$256 after data loading, where $T=1$ for the single image setting and $T>1$ for the video setting. We initialize the ResNet-18 (he2016deep, ) network with the ImageNet-1K (deng2009imagenet, ) pre-trained weights for the multi-scale features extraction. The ResNet-18 (he2016deep, ) backbone returns feature at five levels, with feature sizes as follows given the thermal images as input:
[…]
${Features([64,160,128],[64,80,64],[128,40,32],}$
[…]
We chose the third level features for future processing considering they carry fine-grained spatial details, which is crucial for accurate mask generation and spatial reasoning in our spatial transformer.
Mask-guided Spatial Transformer Module. For spatial feature refinement, we implement a mask-guided attention mechanism that emphasizes regions of interest within the thermal images, specifically focusing on the hands’ positions. This approach utilizes predicted masks from a segmentation head with two convolution layers from the mid-level feature to guide the attention mechanism.
The dimension of the spatial self-attention module we used is 1024, for which we use an additional one-stride convolution to lift the number of channels of the original mid-level feature. Two transformer encoder layers are leveraged and the transformer head number $M$ is set to be 8. The deformable attention mechanism introduced in deformable DETR (zhu2020deformable, ) is used and the number of sampling points per attention head $K$ is 4:
(8) $\small\text{DA}(z_{q},p_{q},x)=\sum_{m=1}^{M}W_{m}\left[\sum_{k=1}^{K}A_{mqk}\cdot W^{\prime}_{m}\left(x(p_{q}+\Delta p_{mqk})\right)\right]$
where $x$ is the masked low-level feature, $m$ indexes the attention head, $\Delta p_{mqk}$ and $A_{mqk}$ denote the sampling offset and the learnable attention weight for the $k$-th sampling point in the $m$-th attention head respectively. Particularly, $A_{mqk}$ is a scalar attention weight lying between 0 and 1, and normalized such that the sum across all $K$ points is 1, $\Delta p_{mqk}$ is a pair of 2D real numbers representing an unconstrained range. Noted that 2-D reference point $p_{q}$ is selected from the hand area derived from the predicted mask and $\Delta p_{mqk}$ and $A_{mqk}$ are obtained via linear projection over the query feature $z_{q}$. In our self-attention, $z_{q}$ is the added feature for $x$ and the corresponding position encoding. Following deformable DETR (zhu2020deformable, ), the first $2M\times K$ channels encode the sampling offsets and the remaining $M\times K$ channels are for the softmax operation to obtain the attention weights.
The output spatial embedding of the deformable mask-guided spatial attention module has the same size as the input feature, which is [1024, 40, 32]. Three convolution layers with a stride of 2 are used to downsample the feature map to [128, 10, 8].
Temporal Transformer Module. This module is crucial for understanding temporal patterns over consecutive frames, utilizing a transformer-based approach. Given per-frame downsample feature map from the spatial transformer module, we flattened them and apply a linear layer to project this spatial embedding to [1,512]. Per-frame spatial feature vectors are then fed to the temporal self-attention, which attends to the feature vector of every frame. The number of attention layers is $2$ and eight heads are used for each attention. The resulting spatio-temporal embedding has the same dimension as the input spatial features, which is [$T$,512].
Pose Regression. In the pose head, we use the MLP to project the spatio-temporal embedding to the output space and obtain the per-frame 3D joint $\mathcal{J}_{i}$ for each frame individually.
$MLP(512\rightarrow 512\rightarrow 512\rightarrow 42\times 3)$
Loss Functions. Two losses are used for joint learning of hand segmentation and pose estimation during the training process. The L1 hand pose regression loss is defined as:
(9) $\footnotesize L_{Hand}=\|P^{2D}-P^{2D}_{gt}\|_{1}+\lambda_{1}\|P^{depth}-P^{depth}_{gt}\|_{1}+\lambda_{2}\|P^{3D}-P^{3D}_{gt}\|_{1}$
where $P^{3D}$, $P^{3D}_{gt}$,$P^{2D}$, $P^{2D}_{gt}$ are the 3D/2D projection of estimated hand joint positions and ground truth hand joint positions. $P^{depth}$ and $P^{depth}_{gt}$ are the corresponding depth values for the hand joints. $\lambda_{1}$ and $\lambda_{2}$ are the weight parameters set as 100 in our implementation. The hand mask binary cross entropy loss is formulated as follows:
(10) $L_{Mask}$ $=-\frac{1}{WH}\sum_{w=1}^{W}\sum_{h=1}^{H}[w_{pos}\cdot M_{wh}\cdot\log(\hat{M}_{wh})$
$+w_{neg}\cdot(1-M_{wh})\cdot\log(1-\hat{M}_{wh})]$
where $M_{wh}$ is the binary ground truth label at pixel location $(w,h)$ and $\hat{M}_{wh}$ is the predicted probability for the estimated mask. $w_{pos}$ and $w_{neg}$ are the weights to cope with the foreground-background imbalance. We set them to $30:1$ during our training.
Training. We apply the Adam optimizer for our baseline training with an initial learning rate of 1e-4, decayed by a coefficient of 0.95 every epoch. All experiments are conducted using two RTX 3090 GPUs, with a batch size of 16 per GPU. Models are trained for 25 epochs.